a triangle sides 2y, 6y, and y + 7. what is the perimeter of the triangle when y=3?
Answer by richard1234(7193)

The problem is flawed. Normally, the perimeter would be 2y + 6y + (y+7) = 9y + 7 = 27 + 7 = 34, but if you substitute y first, you obtain a "triangle" with sides 6, 18, and 10. This is not a triangle, because 6 + 10 < 18, which violates the triangle inequality.